x+y=56 y=3x solve these using substitution, The number 56 equals the number of houses available for sale, and there are two different floor plans available (x&y)I am trying to determine how many of each type there is.
Answer by stanbon(75887)

x+y=56 y=3x solve these using substitution,
---
Substitute for "y" and solve for "x":
x + 3x = 56
4x = 56
x = 14
---
Since y = 3x, y = 3*14 = 42
==================================
Cheers,
Stan H.